a 10kg stone is suspended with a rope of breaking strength 30 kg-wt. the min time in which the stone can be raised through a height 10m starting from rest is ( taking g =10N/kg)
(here what exactly is the breaking stress of the rope)

Grade:Upto college level

5 Answers

Vignesh M
9 years ago
Breaking Stress Is 30 kg-wt = 300N a = f/m = 300/10 = 30m/s Net acceleration = 30 = g + a threfore a = 20m/s s = ut + 0.5at*t threfore 10 = 0 + 0.5*20*t*t 10 = 10*t*t t = 1s. If happy = Approve me
vansh kharbanda
9 years ago
but can u also explain what is breaking stress ??
Vignesh M
9 years ago
lllfescsfs
Vignesh M
9 years ago
Stress Is Actually Force Per Unit Area. But Here Breaking Stress means to maximum Force which the string can sustain. If Breaking stress in 300 N then even 300.001 N would break the string. For calculation purposes just use the value of Breaking stress as Max. Force or Just Force.
vansh kharbanda
9 years ago
thank you.